OpenStudy (rainbow_rocks03):

PLZ help will MEDAL+FAN for BEST response. The table shows the results of drawing one marble from a bag of colored marbles. Each marble was replaced after the outcome was recorded. If there are a total of 50 marbles in the bag, about how many of the marbles are orange? Color Number of Pulls red 5 green 6 orange 2 black 7 A. 5 B. 20 C. 2 D. 10

OpenStudy (rainbow_rocks03):

@kropot72

OpenStudy (rainbow_rocks03):

No direct answers give out the explanation.

OpenStudy (rainbow_rocks03):

@TheSmartOne

OpenStudy (rainbow_rocks03):

@NeonStrawsForever

OpenStudy (hick4life):

Here you go this is your answer 6 + 4 + 7 + 3 = 20 Marbles per pull. Total marbles = 50. 50 ÷ 20 = 2.5 4 * 2.5 = 10 green marbles expected

OpenStudy (anonymous):

@hick4life Direct answers aren't allowed.

OpenStudy (rainbow_rocks03):

I will tag yall if I need more help

OpenStudy (hick4life):

owch